OpenStudy (anonymous):

The function, f(x) = x² - bx - 2b (where b is a positive constant), after completing the square... f(x) = ( x - (b/2) )² - 2b - (b²/4). can i know why is (b/2) the line of symmetry of the graph ? how does this come about ? The minima of the graph becomes ( b/2, - 2b - (b²/4) )

OpenStudy (anonymous):

swag

OpenStudy (anonymous):

You understand how to get to f(x) = )x-(b/2))^2 -(b^2/4) , right? Just the symmetry part?

OpenStudy (anonymous):

yes jus the symmetry part, i only know y = mx + c @.@

OpenStudy (anonymous):

im confuse about the mathematical law/rule that from Completing the square we can find the coordinates of maximum/minimum O:

OpenStudy (anonymous):

have you learned about conics yet? like this equation y= a(x-h)^2 +k

OpenStudy (anonymous):

hmm.. not sure o.o

OpenStudy (anonymous):

this is a general picture of axis of symmetry. Definition is that axis of symmetry divides the parabola into two perfect, identical havles. http://www.mathwarehouse.com/geometry/parabola/images/picture-of-axis-of-symmetry-thumb.png lemme know if anything confuses you

OpenStudy (anonymous):

uh ok ty :)

OpenStudy (anonymous):

axis of symmetry is always a straight, vertical line. let's start with the easiest example: y= x^2 http://www.bbc.co.uk/schools/gcsebitesize/maths/images/graph_24.gif by looking at the graph, can you tell me what's the equation for the axis of symmetry? x=?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

good, :) now are you familiar with translations?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

let's try this one for example. find axis of symmetry of the blue line y= (x-2)^2 http://900igr.net/datai/algebra/Grafik-funktsii-Y-X/0004-004-Takim-obrazom-grafik-funktsii-yx-22-mozhet-byt-poluchen-iz.png

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

oh I should have covered up the answer xDD

OpenStudy (anonymous):

oic! i think i get it

OpenStudy (anonymous):

so it comes from translation

OpenStudy (anonymous):

thank you study100 xD

OpenStudy (anonymous):

Yes, since y= (x-2)^2 is a graph of x^2 that shifts right 2 units, the line of symmetry also shifts 2 units. And yes it does involve translation :)

OpenStudy (anonymous):

if you have time, I suggest you starting to learn conics section: parabola. It will make completing the squares/ finding maximum /minimum and such 100x easier. These are very helpful tutorials , watch from1 to 4 https://www.khanacademy.org/math/algebra2/conics_precalc/parabolas_precalc/v/parabola-intuition-example-1 feel free to lemme know whenever you need help with this :)

OpenStudy (anonymous):

btw, do you know how to find vertex? vertex is the lowest or highest point in a parabola, which is max/min of the quadratic equation.

OpenStudy (anonymous):

uhm.. sorry i dont know =.=

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

this site has the perfect way of explaining parabolas. This is all you really need to know about vertex http://www.mathwarehouse.com/geometry/parabola/vertex-of-a-parabola.php

OpenStudy (anonymous):

by using the formula to find axis of symmetry x= b/2a use that x, plug x in the orginal ax^2 + bx + c (for ex) and find y. then that order pair (x,y) is your vertex.

OpenStudy (anonymous):

plug the x of axis symmetry ***